OpenStudy (anonymous):

For f(x)=3x+1and g(x)=x^2-6, find (fog)(4)

OpenStudy (zale101):

fog is similar to f(g(x)) now, first, apply x=4 into the function of g and use the result to replace with the function of f f(x^2-6) f(4^2-6) f(10) now apply 10 to the function of f

OpenStudy (anonymous):

wait what?

OpenStudy (anonymous):

@Zale101

OpenStudy (zale101):

f(g(x)) is another way of saying f composition to g f(g(x)) and we know that x = 4 right? so f(g(4)) what's function g g(x)=x^2-6 so plug that into the composition and since the function of g(x) is g(4), that means that the x-values of g is 4, correct? f(g(x^2-6)) f(g(4^2-6)) f(16-6) f(10) so the function of f has an x value of 10 so plug the x value 10 to the function of f f(x)=3x+1 (3(10)+1) ?

OpenStudy (the_fizicx99):

Plug in 4 for x into g(x) \(\ g(x) \rightarrow g(4) \) \(\ g(4) = (4)^2 - 6 \rightarrow 16 - 6 = 10 \) \(\ g(4) = 10 \) Now plug that 10 into f(x) \(\ f(10) = 3(10) + 1 \)

OpenStudy (zale101):

thanks @tHe_FiZiCx99

OpenStudy (zale101):

the letex was good

OpenStudy (the_fizicx99):

;) you rock @Zale101

OpenStudy (zale101):

:3

OpenStudy (the_fizicx99):

;>

OpenStudy (anonymous):

so you'd get 31?

OpenStudy (zale101):

yes

OpenStudy (anonymous):

oh okay, thanks a lot! that really helped!

OpenStudy (zale101):

glad it did!

OpenStudy (the_fizicx99):

Lol you're welcome ;) ... even though I'm late and didn't see your reply >.>